# Variational inequalities and smooth-fit principle for singular stochastic control problems in Hilbert spaces

### v1

@@ Abstract.
We consider a class of infinite-dimensional singular stochastic control problems. These can be thought of as spatial monotone follower problems and find applications in spatial models of production and climate transition. Let $(D,\mathcal{M},\mu)$ be a finite measure space and consider the Hilbert space $H:=L^{2}(D,\mathcal{M},\mu;\mathbb{R})$. Let then $X$ be an $H$-valued stochastic process on a suitable complete probability space, whose evolution is determined through an SPDE driven by a self-adjoint linear operator $\mathcal{A}$ and affected by a cylindrical Brownian motion. The evolution of $X$ is controlled linearly via an $H$-valued control consisting of the direction and the intensity of action, a real-valued nondecreasing right-continuous stochastic process, adapted to the underlying filtration. The goal is to minimize a discounted convex cost-functional over an infinite time-horizon. By combining properties of semiconcave functions and techniques from viscosity theory, we first show that the value function of the problem $V$ is a $C^{1,Lip}(H)$-viscosity solution to the corresponding dynamic programming equation, which here takes the form of a variational inequality with gradient constraint. Then, by allowing the decision maker to choose only the intensity of the control and requiring that the given control direction $\hat{n}$ is an eigenvector of the linear operator $\mathcal{A}$, we establish that the directional derivative $V_{\hat{n}}$ is of class $C^{1}(H)$, hence a second-order smooth-fit principle in the controlled direction holds for $V$. This result is obtained by exploiting a connection to optimal stopping and combining results and techniques from convex analysis and viscosity theory.
@@ 1. Introduction
Let us describe the class of infinite-dimensional singular stochastic control problems and our contributions more precisely. Let $(D,\mathcal{M},\mu)$ be a finite measure space, and consider the Hilbert space $H:=L^{2}(D,\mathcal{M},\mu;\mathbb{R})$. The state variable is described by a stochastic process $X$ with values in $H$. Its evolution is determined by a stochastic partial differential equation (SPDE) driven by a self-adjoint linear operator $\mathcal{A}$ and a cylindrical Brownian motion. Next to technical requirements, we assume that the operator $\mathcal{A}$ generates a $C_{0}$-semigroup of positivity-preserving contractions. As a benchmark case, one can consider the sum of the Laplacian operator and a multiplicative operator of the form $-\delta x$, for $\delta>0$, representing a depreciation or dissipative term. The evolution of $X$ is controlled linearly through an $H$-valued control, encompassing both direction and intensity of action, the latter being a real-valued nondecreasing right-continuous stochastic process adapted to the underlying filtration. The objective is to minimize a discounted convex cost-functional of the form
@@
We proceed by deriving the dynamic programming equation associated with (1.1) and demonstrate that $V$ is a $C^{1,\text{Lip}}(H)$-viscosity solution to it. The proof of this result relies on a Dynkin’s formula for test functions of semimartingales and utilizes an equivalent representation of $V$, derived from a tailored application of the Radon-Nikodym theorem for vector-valued measures (see Lemma 2.11 and (2.19) below). It is noteworthy that the proof of the supersolution property of $V$ employs a novel argument derived from an inequality stemming from Dynkin’s formula and the dissipativity property of the operator $\mathcal{A}$. This approach is applicable in finite-dimensional settings as well, substantially reducing the technicalities usually associated with demonstrating the supersolution (or subsolution) property in minimization (or maximization) problems involving singular controls (see, e.g., [16], [34, Ch. VIII], [38], and [49], among others).
To further enhance the regularity of $V$, we introduce the assumption that the decision-maker can only control the intensity of action. The direction of action, denoted by $\hat{n}\in H$, is then taken to be an eigenvector of the operator $\mathcal{A}$. Under this requirement and further technical properties of $G$, we are able to show that the directional derivative of $V$ in the direction of control, $V_{\hat{n}}$, is such that $V_{\hat{n}}\in C^{1}(H)$. This result can be read as a second-order smooth fit property of $V$, a regularity result of particular relevance in singular stochastic control problems (see the discussion in Section 5.1 below). The aforementioned smooth-fit property is obtained by identifying $V_{\hat{n}}$ as the value function of an optimal stopping problem (in the spirit of the finite-finite dimensional contribution [4]) and subsequently examining the regularity of its (sub)gradient. In particular, under a suitable nondegeneracy condition on the Brownian noise, assuming that the directional derivative $G_{\hat{n}}$ is semiconcave, and combining arguments from viscosity theory and convex analysis, we are able to show that $V_{\hat{n}}$ is Fréchet differentiable at any $x\in H$ and that the gradient $DV_{\hat{n}}\in C(H;\mathcal{D}(\mathcal{A}))$ (with the domain $\mathcal{D}(\mathcal{A})$ being endowed with the graph norm). For further details, please refer to Proposition 4.17.
@@
The theory of regular stochastic control and of optimal stopping in infinite-dimensional (notably, Hilbert) spaces received a large attention in the last decades (see, e.g., the monography [27] for control problems, and [5], [18], [19], [29], [35], [36], [59] for optimal stopping). As previously discussed, we contribute to that bunch of literature by providing the viscosity property and $C^{1}$-regularity (smooth-fit) of the value function of a class of optimal stopping problems in Hilbert spaces. To the best of our knowledge, such a regularity result appears here for the first time, and it is therefore of independent interest.
@@
Organization of the Paper. The rest of the paper is organized as follows. In Section 2 we provide the setting and introduce the problem. In Section 3 we then consider the variational inequality associated to the problem and prove preliminary regularity and viscosity property of its value function $V$. Under a suitable requirement on the direction of action, in Section 4 a connection to optimal stopping is derived and regularity of the optimal stopping problem’s value function is proved. As a byproduct of that, in Section 5 a second-order smooth-fit property for $V$ is then obtained. Finally, Section 6 proposes two applications in Economics, while Appendix A collects a result on semiconcave and semiconvex functions and Appendix B technical lemmata.
@@ 2.1. Setting
In particular (see, e.g., [26, Ch. II, Sec. 3] and [8, Ch. II-1, Sec. 2.10.1]), under Assumption 2.1, the operator $\mathcal{A}$ generates a $C_{0}$-semigroup of contractions $(e^{t\mathcal{A}})_{t\geq 0}\subseteq\mathcal{L}(H)$ and
@@
Furthermore, we impose the next assumption.
Assumption 2.4.
$\mathcal{D}(\mathcal{A})\hookrightarrow L^{\infty}(D,\mathcal{M},\mu;\mathbb{R}).$
Remark 2.5.
In the examples considered in Section 6, $\mathcal{D}(\mathcal{A})$ will be the Sobolev space $W^{2,2}(\mathcal{O})$ with appropriate boundary conditions, for $\mathcal{O}$ being an open, simply connected, and bounded set of $\mathbb{R}^{n}$ ($n<4$) with smooth boundary. Assumption 2.4 is then verified in this setting by [12, Cor. 9.15].
Let us now come to the probabilistic structure of our setup. We endow the time-interval $[0,\infty)$ with the Borel $\sigma$-algebra $\mathcal{B}([0,\infty))$. Also, let $(\Omega,\mathcal{F},\mathbb{F},\mathsf{P})$ be a filtered probability space, with filtration $\mathbb{F}:=(\mathcal{F}_{t})_{t\in[0,\infty)}$ satisfying the usual conditions, and let $W$ be a cylindrical Wiener process on $(\Omega,\mathcal{F},\mathbb{F},\mathsf{P})$, taking values in another Hilbert space $K$. Finally, for future use, we denote by $\mathcal{T}$ the set of all $\mathbb{F}$-stopping times.
@@
(2.1) $\mathcal{M}$ ::=:= $\big{\{}I:\Omega\times[0,\infty)\to H_{+}:\,\text{$I_{\cdot}$ is}\,\,\mathbb{F}-\text{adapted and such that}\,\,t\mapsto I_{t}$
@@
By setting $I_{0^{-}}:=0$ for any $I\in\mathcal{M}$, we mean that we extend any $I\in\mathcal{M}$ by setting $I\equiv 0$ on $[-\varepsilon,0)$, for a given and fixed $\varepsilon>0$. In this way, the associated measures have a positive mass at initial time of size $I_{0}$. Notice that this is equivalent with identifying any control $I$ with a countably additive measure $I:\mathcal{B}([0,\infty))\to[0,\infty)$ of local finite variation defined as $I((s,t]):=I_{t}-I_{s}$, for every $s,t\in[0,\infty)$, $s<t$, plus a Dirac-delta at time 00 of amplitude $I_{0}$.
@@
Denoting by $\mathcal{L}_{1}(H)$ the set of nonnegative trace-class operators on $H$ and endowing $\mathcal{L}_{1}(H)$ with the usual norm
$|Q|_{\mathcal{L}_{1}(H)}:=\mbox{Tr}[Q]=\sum_{k=0}^{\infty}\langle Qe_{k},e_{k}\rangle_{H},$
where $(e_{k})$ is any orthonormal basis of $H$, we then have under Assumption 2.7 that
$\sigma\sigma^{*}\in\mathcal{L}_{1}(H).$
Notice that Assumptions 2.1 and 2.7 imply that the stochastic convolution (2.9) is well defined and continuous (see [21, Ch. 5]). For future use, we also note that, because of Assumptions 2.1 and 2.7, for all $m\in[1,\infty)$ one has for some $\overline{c}_{m}>0$ (see [37])
@@
(2.11) $\mathsf{E}\left[\sup_{t\geq 0}|X^{x,0}_{t}|^{m}_{H}\right]\leq c_{p}(1+|x|^{m}_{H}),\ \ \ \ \forall x\in H,$
@@ 2.2. Problem formulation
We now move on by introducing the infinite-dimensional singular stochastic control problem which is the object of our study. Let
$G:H\to\mathbb{R}$
be a running cost function, satisfying the following requirements.
@@
    ${G}$ is convex; There exists $c_{o},\kappa_{1},\kappa_{2}>0$ and $p\geq 2$ such that
    $\kappa_{1}|x|_{H}^{p}-\kappa_{2}\leq{G}(x)\leq c_{o}(1+|x|_{H}^{p});$
@@
    $G$ is semiconcave with semiconcavity constant $c_{1}>0$; that is, there exists $c_{1}>0$ such that
    $\lambda G(x)+(1-\lambda)G(y)-G(\lambda x+(1-\lambda)y)\leq\frac{c_{1}}{2}\lambda(1-\lambda)|x-y|_{H}^{2},\ \ \forall x,y\in H,\ \lambda\in[0,1].$
Notice that by Lemma (i)(ii) one has that $G\in C^{1,\text{Lip}}(H)$.
@@
Benchmark examples satisfying Assumption 2.8 are the quadratic cost function
$G(x)=\frac{1}{2}|x-\overline{x}|^{2}_{H},\ \ \ \ x\in H,$
for some target level $\overline{x}\in H$, as well as
${G}(x)=\frac{1}{2}\langle x,h\rangle_{H}^{2},\quad\text{or}\quad{G}(x)=\frac{1}{2}\langle Qx,x\rangle_{H},\quad x\in H,$
with $h\in H$, and with $Q$ being positive semidefinite and symmetric.
Let now (cf. (2.1))
(2.12) $\mathcal{I}$ ::=:= $\Big{\{}I\in\mathcal{M}:\,\,\,\mathsf{E}\bigg{[}\Big{|}\int_{0^{-}}^{T}|\hat{\vartheta}_{s}|_{H}\,\mathrm{d}|I|_{s}\Big{|}^{p}\bigg{]}<\infty\,\,\,\text{for}\,\,p\geq 2\,\,\text{as in}$
[…]
be the class of admissible controls. For a discount rate $\rho>0$ and for $q\in H_{+}$ such that $q\geq q_{o}\mathbf{1}$ for some $q_{o}>0$ (being $\mathbf{1}\in H$ the constant unitary vector of $H$), recalling (2.1) we introduce the expected cost functional
which is well-defined, although potentially infinite. The infinite-dimensional singular stochastic control problem under study is then
@@
(2.16) $\mathcal{S}$ ::=:= $\{\nu:\Omega\times[0,\infty)\to[0,\infty):\,\text{$\nu_{\cdot}$ is}\,\,\mathbb{F}-\text{adapted and such that}\,\,t\mapsto\nu_{t}$
@@
(2.17) $\mathcal{I}_{0}$ ::=:= $\Big{\{}(\vartheta,\nu):\Omega\times[0,\infty)\to\Theta\times[0,\infty):\ \vartheta_{\cdot}\ \mbox{is}\ \mathbb{F}-\mbox{adapted},\,\,\nu_{\cdot}\in\mathcal{S}\,\,\,\text{and}$
[…]
@@
For $p\geq 2$ as in Assumption 2.8 (see also (2.12)), one clearly has that the integrability conditions required in (2.17) are met, because of the previous definitions of $\vartheta$ and $\mathrm{d}\nu$, and because $I\in\mathcal{I}$. This shows the first part of the claim.
@@
In the next Section 3, we will make use of both the equivalent representations (2.14) and (2.20). In particular, preliminary regularity properties of $V$ (see Section 3.1 below) as well as Propositions 3.3 and 3.4 in Section 3.2 will be shown by using (2.14), while the viscosity property of $V$ will be proved through (2.20) (see Proposition 3.5 and Theorem 3.7 in Section 3.2).

### v2

@@ Abstract.
We consider a class of infinite-dimensional singular stochastic control problems. These can be thought of as spatial monotone follower problems and find applications in spatial models of production and climate transition. Let $(D,\mathcal{M},\mu)$ be a finite measure space and consider the Hilbert space $H:=L^{2}(D,\mathcal{M},\mu;\mathbb{R})$. Let then $X$ be an $H$-valued stochastic process on a suitable complete probability space, whose evolution is determined through an SPDE driven by a self-adjoint linear operator $\mathcal{A}$ and affected by a cylindrical Brownian motion. The evolution of $X$ is controlled linearly via an $H$-valued control consisting of the direction and the intensity of action, a real-valued nondecreasing right-continuous stochastic process, adapted to the underlying filtration. The goal is to minimize a discounted convex cost-functional over an infinite time-horizon. By combining properties of semiconcave functions and techniques from viscosity theory, we first show that the value function of the problem $V$ is a $C^{1,\mathrm{Lip}}(H)$-viscosity solution to the corresponding dynamic programming equation, which here takes the form of a variational inequality with gradient constraint. Then, by allowing the decision maker to choose only the intensity of the control and requiring that the given control direction $\hat{n}$ is an eigenvector of the linear operator $\mathcal{A}$, we establish that the directional derivative $V_{\hat{n}}$ is of class $C^{1}(H)$, hence a second-order smooth-fit principle in the controlled direction holds for $V$. This result is obtained by exploiting a connection to optimal stopping and combining results and techniques from convex analysis and viscosity theory.
@@ 1. Introduction
Let us describe the class of infinite-dimensional singular stochastic control problems and our contributions more precisely. Let $(D,\mathcal{M},\mu)$ be a finite measure space, and consider the Hilbert space $H:=L^{2}(D,\mathcal{M},\mu;\mathbb{R})$. The state variable is described by a stochastic process $X$ with values in $H$. Its evolution is determined by a stochastic partial differential equation (SPDE) driven by a self-adjoint linear operator $\mathcal{A}$ and a cylindrical Brownian motion. Next to technical requirements, we assume that the operator $\mathcal{A}$ generates a $C_{0}$-semigroup of positivity-preserving contractions. As a benchmark case, one can consider the sum of the Laplacian operator and a multiplicative operator of the form $\delta x$, for $\delta>0$, representing a depreciation or dissipative term. The evolution of $X$ is controlled linearly through an $H$-valued control, encompassing both direction and intensity of action, the latter being a real-valued nondecreasing right-continuous stochastic process adapted to the underlying filtration. The objective is to minimize a discounted convex cost-functional of the form
@@
We begin by deriving the dynamic programming equation associated with (1.1) and show that $V$ is a $C^{1,\text{Lip}}(H)$-viscosity solution of this equation. The proof relies on Dynkin’s formula for test functions of semimartingales and on an alternative representation of $V$, obtained through a tailored application of the Radon–Nikodym theorem for vector-valued measures (see Lemma 2.10 and (2.21) below). In particular, the proof of the supersolution property of $V$ introduces a new argument based on an inequality derived from Dynkin’s formula together with the dissipativity property of the operator $\mathcal{A}$. This approach also extends naturally to finite-dimensional settings, greatly simplifying the usual technical difficulties in proving the supersolution (or subsolution) property in minimization (or maximization) problems involving singular controls (see, e.g., [16], [35, Ch. VIII], [39], and [51]).
To further enhance the regularity of $V$, we introduce the assumption that the decision-maker can only control the intensity of action. The direction of action, denoted by $\hat{n}\in H$, is then taken to be an eigenvector of the operator $\mathcal{A}$. Under this requirement and further technical properties of $G$, we are able to show that the directional derivative of $V$ in the direction of control, $V_{\hat{n}}$, is such that $V_{\hat{n}}\in C^{1}(H)$. This result can be read as a second-order smooth fit property of $V$, a regularity result of particular relevance in singular stochastic control problems (see the discussion in Section 5.1 below). The aforementioned smooth-fit property is obtained by identifying $V_{\hat{n}}$ as the value function of an optimal stopping problem (in the spirit of the finite-finite dimensional contribution [4]) and subsequently examining the regularity of its (sub)gradient. In particular, under a suitable nondegeneracy condition on the Brownian noise, assuming that the directional derivative $G_{\hat{n}}$ is semiconcave, and combining arguments from viscosity theory and convex analysis, we are able to show that $V_{\hat{n}}$ is Fréchet differentiable at any $x\in H$ and that the gradient $DV_{\hat{n}}\in C(H;H)$. For further details, please refer to Proposition 4.19.
@@
The theory of regular stochastic control and of optimal stopping in infinite-dimensional (notably, Hilbert) spaces received a large attention in the last decades (see, e.g., the monography [28] and the recent regularity results in [23] for control problems, and [5], [18], [19], [30], [36], [37], [63], [64] for optimal stopping). As previously discussed, we contribute to that bunch of literature by providing the viscosity property and $C^{1}$-regularity (smooth-fit) of the value function of a class of optimal stopping problems in Hilbert spaces. To the best of our knowledge, such a regularity result appears here for the first time, and it is therefore of independent interest.
@@
Organization of the Paper. The rest of the paper is organized as follows. In Section 2 we provide the setting and introduce the problem. In Section 3 we then consider the variational inequality associated to the problem and prove preliminary regularity and viscosity property of its value function $V$. Under a suitable requirement on the direction of action, in Section 4 a connection to optimal stopping is derived and regularity of the optimal stopping problem’s value function is proved. As a byproduct of that, in Section 5 a second-order smooth-fit property for $V$ is then obtained. Finally, Section 6 proposes two applications in Economics, while Appendix A provides an extended sketch of the proof of the Dynamic Programming Principle and Appendix B some technical lemmata.
@@ 2.1. Setting
In particular (see, e.g., [27, Ch. II, Sec. 3] and [8, Ch. II-1, Sec. 2.10.1]), under Assumption 2.1, the operator $\mathcal{A}$ generates a $C_{0}$-semigroup of contractions $(e^{t\mathcal{A}})_{t\geq 0}\subseteq\mathcal{L}(H)$ and
@@
Let us now come to the probabilistic structure of our setup. We endow the time-interval $0,\infty)$ with the Borel $\sigma$-algebra $\mathcal{B}([0,\infty))$. Also, let $(\Omega,\mathcal{F},\mathbb{F},\mathsf{P})$ be a filtered probability space, with filtration $\mathbb{F}:=(\mathcal{F}_{t})_{t\in[0,\infty)}$ satisfying the usual conditions, and let $W$ be a cylindrical Wiener process on $(\Omega,\mathcal{F},\mathbb{F},\mathsf{P})$, taking values in another Hilbert space $K$. Finally, for future use, we denote by $\mathcal{T}$ the set of all $\mathbb{F}$-stopping times.
@@
(2.1) $\mathcal{M}$ $:=$ $\big\{I:\Omega\times[0,\infty)\to H_{+}:\,\text{$I_{\cdot}$ is}\,\,\mathbb{F}-\text{adapted and such that}\,\,t\mapsto I_{t}$
@@
By setting $I_{0^{-}}:=0$ for any $I\in\mathcal{M}$, we mean that we extend any $I\in\mathcal{M}$ by setting $I\equiv 0$ on $\varepsilon,0)$, for a given and fixed $\varepsilon>0$. In this way, the associated measures have a positive mass at initial time of size $I_{0}$. Notice that this is equivalent with identifying any control $I$ with a countably additive measure $I:\mathcal{B}([0,\infty))\to[0,\infty)$ of local finite variation defined as $I((s,t]):=I_{t}-I_{s}$, for every $\in[0,\infty), s<ts<t$, plus a Dirac-delta at time 0 of amplitude $I_{0}$.
@@
We denote by $\mathcal{L}_{1}(H)$ the set of nuclear operators on $H$, i.e. operators $Q\in\mathcal{L}(H)$ such that
(2.10) $|Q|_{\mathcal{L}_{1}(H)}:=\mbox{Tr}[|Q|]:=\sum_{k=0}^{\infty}\langle\sqrt{Q^{*}Q}e_{k},e_{k}\rangle_{H}<\infty,$
where $(e_{k})$ is any orthonormal basis of $H$. As well known, the above quantity does not depend on the choice of the basis and $\mathcal{L}_{1}$ is a Banach space endowed by the norm defined in (2.10) . We then have, under Assumption 2.5, that $\sigma\sigma^{*}\in\mathcal{L}_{1}(H).$
Notice that Assumptions 2.1 and 2.5 imply that the stochastic convolution (2.9) is well defined and continuous (see [21, Ch. 5]). For future use, we also note that, because of Assumptions 2.1 and 2.5, for all $m\in[1,\infty)$ one has for some $\overline{c}_{m}>0$ (see [38])
@@
(2.12) $\mathsf{E}\left[\sup_{t\geq 0}|X^{x,0}_{t}|^{m}_{H}\right]\leq c_{m}(1+|x|^{m}_{H}),\ \ \ \ \forall x\in H,$
@@
Remark 2.6.
The Ornstein–Uhlenbeck setting plays a fundamental role in guaranteeing – in a direct but still not immediate way – the well-posedness of the state equation in the infinite-dimensional context, thereby providing a natural starting point for the analysis of the control problem. Its linear structure makes it possible to rigorously define the mild solution of the underlying controlled stochastic evolution equation and to derive the necessary a priori estimates on the state process. Moreover, this framework allows for the explicit derivation of gradient estimates and facilitates the application of semi-concavity and regularization techniques within a Hilbert space setting.
Extending the analysis to the relevant case of nonlinear dynamics introduces substantial issues even at the preliminary stage of studying the well-posedness of the controlled state equation, due to the presence of the singular control. We leave a detailed study of this interesting class of problems for future research.
@@ 2.2. Problem formulation
We now move on by introducing the infinite-dimensional singular stochastic control problem which is the object of our study. Let $G:H\to\mathbb{R}$ be a running cost function, satisfying the following requirements.
@@
    G{G} is convex and there exist $c_{o},\kappa_{1},\kappa_{2}>0$ such that κ_1 —x—_H^2 - κ_2 G(x)c_o(1+—x—_H^2);
@@
    $G$ is semiconcave with semiconcavity constant $c_{1}>0$; that is, there exists $c_{1}>0$ such that λG(x)+(1-λ) G(y) - G(λx+(1-λ) y)c12λ(1-λ)—x-y—_H^2, x,yH, λ[0,1].
Notice that by [48] (see end of Page 265 therein) one has that $G\in C^{1,\text{Lip}}(H)$.
@@
  1. (i)
    Benchmark examples satisfying Assumption 2.7 are the quadratic cost function G(x)=12—x-x—^2_H, xH, for some target level $\overline{x}\in H$, as well as G(x)=12x,h_H^2,  or  G(x)=12Q x, x_H,  x H, with $h\in H$, and with $Q$ being positive semidefinite and symmetric.
  2. (ii)
    Assumption 2.7(ii) amounts to requiring semiconcavity with a linear modulus. One could weaken this condition by assuming that, for some $\alpha\in(0,1]$,
    (2.13) $\lambda G(x)+(1-\lambda)G(y)-G(\lambda x+(1-\lambda)y)\leq\frac{c_{1}}{2}\lambda(1-\lambda)|x-y|_{H}^{1+\alpha},$
    for all $x,y\in H$ and for all $\lambda\in[0,1]$. Under this weaker assumption, all the results in this paper remain valid, after making the corresponding adjustments. As a matter of fact, by mimicking the arguments developed in the finite-dimensional setting in [14, Th. 3.3.7], one can still prove that a convex function satisfying (2.13) belongs to the class $C^{1,\alpha}_{loc}(H)$.
For a discount rate $\rho>0$ and for $q\in H_{+}$ such that $q\geq q_{o}\mathbf{1}$ for some $q_{o}>0$ (being $\mathbf{1}\in H$ the constant function of $H$ identically equal to $1$), recalling (2.1) we introduce the expected cost functional
(2.14) $\mathcal{J}(x;I):=\mathsf{E}\bigg[\int_{0^{-}}^{\infty}e^{-\rho t}\Big({G}(X_{t}^{x,I})\mathrm{d}t+\langle q,\mathrm{d}I_{t}\rangle_{H}\Big)\bigg],\quad(x,I)\in H\times\mathcal{I}.$
Here, the the class of admissible controls $\mathcal{I}$ is defined as (cf. (2.1))
(2.15) $\mathcal{I}$ $:=$ $\Big\{I\in\mathcal{M}:\,\,\,\mathsf{E}\bigg[\Big|\int_{0^{-}}^{T}|\hat{\vartheta}_{s}|_{H}\,\mathrm{d}|I|_{s}\Big|^{2}\bigg]+\mathsf{E}\bigg[\int_{0^{-}}^{T}\frac{|\hat{\vartheta}_{s}|^{2}_{H}}{\langle q,\hat{\vartheta}_{s}\rangle}\,\mathrm{d}|I|_{s}\bigg]<\infty\,\,\,\forall T>0\Big\}.$
The infinite-dimensional singular stochastic control problem under study is then
@@
(2.18) $\mathcal{S}$ $:=$ $\{\nu:\Omega\times[0,\infty)\to[0,\infty):\,\text{$\nu_{\cdot}$ is}\,\,\mathbb{F}-\text{adapted and such that}\,\,t\mapsto\nu_{t}$
@@
(2.19) $\mathcal{I}_{0}$ $:=$ $\Big\{(\vartheta,\nu):\Omega\times[0,\infty)\to\Theta\times[0,\infty):\ \vartheta_{\cdot}\ \mbox{is}\ \mathbb{F}-\mbox{adapted},\,\,\nu_{\cdot}\in\mathcal{S}\,\,\,\text{and}$
$\mathsf{E}\bigg[\Big|\int_{0^{-}}^{T}|{\vartheta}_{s}|_{H}\,\mathrm{d}\nu_{s}\Big|^{2}\bigg]+\mathsf{E}\bigg[\int_{0^{-}}^{T}|{\vartheta}_{s}|^{2}_{H}\,\mathrm{d}\nu_{s}\bigg]<\infty\,\,\,\text{for}\,\,\forall T>0\Big\}.$
@@
Recalling (2.15), one clearly has that the integrability conditions required in (2.19) are met, because of the previous definitions of $\vartheta$ and $\mathrm{d}\nu$, and because $I\in\mathcal{I}$. This shows the first part of the claim.
@@
In the next Section 3, we will make use of both the equivalent representations (2.16) and (2.22). In particular, preliminary regularity properties of $V$ (see Section 3.1 below) as well as Propositions 3.2 and 3.3 in Section 3.2 will be shown by using (2.16), while the viscosity property of $V$ will be proved through (2.22) (see Proposition 3.5 and Theorem 3.8 in Section 3.2).